40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Tue Nov 17 11:09:43 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
Tue Nov 17 13:39:37 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
Tue Nov 17 15:23:06 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
Tue Nov 17 21:10:30 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
Wed Nov 18 11:27:33 2009, Koji, josephb, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
Wed Nov 18 12:33:30 2009, Koji, josephb, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
Wed Nov 18 16:24:25 2009, pete, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC
|
|
Message ID: 2290
Entry time: Wed Nov 18 11:27:33 2009
In reply to: 2285
Reply to this: 2291
|
Author: |
Koji, josephb |
Type: |
Configuration |
Category: |
SUS |
Subject: |
ETMY suspension conencted to megatron ADC/DAC |
|
|
Quote: |
Koji, Rana
The megatron DAC was temporaly connected to the suspension electronics for the DAC test. We went down to ETMY as we could not excite the mirror.
The DAC is putting correct voltages to the channels. However, the anti imaging filter test output does not show any signal.
This means something wrong is there in the DAC I/F box or the cables to the AI circuit. We will check those things tomorrow.
The ETMY was restored to the usual configuration.
|
It appears the front panel for the DAC board is mis-labeled. Channels 1-8 are in fact 9-16, and 9-16 are the ones labeled 1-8. We have put on new labels to reduce confusion in the future. |